OpenStudy (anonymous):

(5x+4)(9x+9) is a factorization of what trinomial?

OpenStudy (anonymous):

Just distribute it out. (5x)(9x)+9(5x)+9x(4)+9(4) 45x^2+45x+36x+36 45x^2+81x+36

OpenStudy (anonymous):

5x*9x=45x^2 5x*9=45x 4*9x=36x 4*9=36 than add\[45x^2+45x+36x+36\]\[45x^2+81x+36\]